# Shilong Liu 5, Haoyu Dong6, Zihao Lin4, Yixin Wang1

### v1

@@ Abstract
Multi-Modal Large Language Models (MLLMs), despite being successful, exhibit limited generality and often fall short when compared to specialized models. Recently, LLM-based agents have been developed to address these challenges by selecting appropriate specialized models as tools based on user inputs. However, such advancements have not been extensively explored within the medical domain. To bridge this gap, this paper introduces the first agent explicitly designed for the medical field, named Multi-modal Medical Agent (MMedAgent). We curate an instruction-tuning dataset comprising six medical tools solving seven tasks, enabling the agent to choose the most suitable tools for a given task. Comprehensive experiments demonstrate that MMedAgent achieves superior performance across a variety of medical tasks compared to state-of-the-art open-source methods and even the closed-source model, GPT-4o. Furthermore, MMedAgent exhibits efficiency in updating and integrating new medical tools.
@@
Binxu Li1, Tiankai Yan1, Yuanting Pan1, Zhe Xu2,3, Jie Luo3, Ruiyang Ji4, Shilong Liu 5, Haoyu Dong6, Zihao Lin4, Yixin Wang1thanks: Corresponding authors 1Stanford University,  2CUHK, 3Harvard Medical School, 4Virginia Tech, 5Tsinghua University, 6Duke University {andy0207,yxinwang}@stanford.edu, haoyu.dong151@duke.edu, zihaol@vt.edu
@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) have made considerable progress across diverse tasks with inputs from different medical imaging modalities (e.g., Magnetic Resonance Imaging, Computed Tomography, X-ray) in healthcare, including Visual Question Answering (VQA) Moor et al. (2023); Zhang et al. (2023a); Li et al. (2023), image classification Sun et al. (2023), image segmentation Ma et al. (2024a), and Medical Report Generation (MRG) Thawkar et al. (2023); Hamamci et al. (2024), etc. Despite these advancements, MLLMs often exhibit limitations in seamlessly solving multiple tasks across different medical imaging modalities. Although recent large medical models Zhang et al. (2023b); Tu et al. (2024); Wu et al. (2023); Yang et al. (2024); Zhao et al. (2024a) have attempted to address this challenge, they remain limited to handling a narrow range of tasks across a restricted set of imaging modalities and cannot be efficiently extended to new tasks or more imaging modalities. Furthermore, these generalists typically do not provide expert-level responses comparable to those of specialized MLLMs customized for specific tasks.
@@
One way to address this issue is to build an AI Agent, an AI system driven by Large Language Models (LLMs) that integrates various domain expert models as tools. Such a system can understand user instructions, make decisions, and select the appropriate tools to execute any specific task, thereby generating expert-level responses for any given request Xie et al. (2024); Chen et al. (2023); Wang et al. (2024); Liu et al. (2023a); Tao et al. (2023). Despite the significant success of AI agents in the general image domain Tao et al. (2023); Qin et al. (2023); Wang et al. (2023a), no such agents currently exist in the medical domain. Although several works Tang et al. (2023); Schmidgall et al. (2024); Li et al. (2024); Fan et al. (2024) in the medical field use the term “agent” in their methods, they focus on utilizing LLMs to play various roles and collaborate on complex tasks, in which an “agent” refers to a specific role.
@@
In this work, we aim to build the first AI agent specifically for the medical domain, termed as Multi-modal Medical Agent (MMedAgent). We choose LLaVA-Med Li et al. (2023) as the backbone and aim to extend its capability to handle various language and multi-modal tasks, including grounding, segmentation, classification, grounding, MRG, and Retrieval-Augmented Generation (RAG). The first step to building MMedAgent is to collect the state-of-the-art (SOTA) methods for each task, hereafter referred to as “tools”. During this phase, we identify a lack of an effective tool for the grounding task, prompting us to fine-tune Grounding DINO Liu et al. (2023b) specifically for medical applications. Next, we build an instruction-based dataset that teaches the agent to select the proper tool(s) when encountering a user instruction and aggregate the outputs from tools to reply to users precisely and comprehensively. The core of our approach involves an end-to-end training regimen through visual instruction tuning Liu et al. (2023a). MMedAgent has demonstrated promising results in various aspects. When evaluated on several complex medical tasks, MMedAgent significantly outperforms current open-source SOTA methods, LLaVA-Med Li et al. (2023) and RadFM Wu et al. (2023), and even surpasses close-source method, GPT-4o OpenAI (2024), on average. It also enhances the backbone’s, i.e., LLaVA-Med, original capability in the VQA task, as well as exhibits efficient capability in learning new tools.
@@
  • We build the first open-source instruction tuning dataset for multi-modal medical agents.
@@ 2.1 Medical MLLMs
Recent progress on LLMs has been made on multi-modal conversational capability Moor et al. (2023); Zhang et al. (2023b); Tu et al. (2024); Zhang et al. (2023c, a); Thawkar et al. (2023); Sun et al. (2023); Wu et al. (2023); Li et al. (2023); Ma et al. (2024a); Yang et al. (2024); Zhao et al. (2024a); Hamamci et al. (2024). Owing to the diversity inherent in medical data and tasks, LLMs have initially been localized to specific imaging domains such as X-ray Thawkar et al. (2023), CT Hamamci et al. (2024), and histology Sun et al. (2023), or tailored for different tasks such as segmentation Ma et al. (2024a); Lei et al. (2023) and medical report generation Wu et al. (2023). In contrast, generalist models expand these capabilities by enabling a single LLM to cover a wider range of imaging modalities and tasks by enlarging the pre-training datasets greatly Zhang et al. (2023b); Li et al. (2023); Zhao et al. (2024a); Liu et al. (2023a); Yang et al. (2024). Although generalist models are capable of handling a wide range of medical modalities and tasks, they face limitations in scalability when incorporating additional skills and lack specialization in specific tasks.
@@ 2.2 AI Agent
A multi-modal AI Agent is a system that achieves users’ general-purpose goals by perceiving the environment and making decisions based on the perceptions Xie et al. (2024); Wooldridge and Jennings (1995). Recent works utilize LLMs as planners to understand multi-modal input from environments and make decisions to call different tools to achieve goals. Based on whether the LLM is open source or not, (Xie et al., 2024) classifies multi-modal AI Agents into two types: (i) closed-source LLMs as planners, which utilize prompt technique to enable LLMs to make decisions Chen et al. (2023); Wang et al. (2024); (ii) fine-tuned LLMs as planners, where an LLM is fine-tuned to understand instructions, make decisions, and call tools/APIs Liu et al. (2023a); Tao et al. (2023). Our MMedAgent belongs to the second type.
Multi-modal AI Agents have achieved great success in various applications. For example, (Tao et al., 2023; Gur et al., 2023; Zhan and Zhang, 2023) apply agents to control the website or user interface. Some works Qin et al. (2023); Wang et al. (2023c) focus on robotics or embodied AI which applies multi-modal LLMs to perceive and interact with real environments. Most works concentrate on multi-modal understanding, editing, or generation, especially image, video, or audio Liu et al. (2023a); Wang et al. (2023a); Zhang et al. (2023d). However, these works are limited to the natural domains, leaving the applications in the medical domain unexplored, which is particularly challenging due to its diverse modalities and tasks. To the best of our knowledge, we are the first to address this challenge and build a system that integrates these varied medical applications.
@@ 3.1 Workflow
Following LLaVA-Plus Liu et al. (2023a), MMedAgent is built to learn to utilize a wide range of multi-modal medical tools, extending the MLLMs’ capabilities to analyze and accomplish various medical tasks. As shown in Figure 1, the workflow consists of four parts: (1) users provide an instruction $X_{q}$ and a medical image $I_{q}$; (2) MLLM works as an action planner, which understands $X_{q}$ and $I_{q}$ and then generates a formatted instruction $X_{\text{tool}}$ to call a specific tool. (3) The tool is executed given $I_{q}$ and the output $X_{\text{result}}$ of the tool is sent to the MLLM. (4) The MLLM aggregates the output with $X_{q}$ and $I_{q}$ and generates the final answer $X_{\text{answer}}$ to users. We train the agent end-to-end with an auto-regressive objective on the generated sequence - $X_{\text{tool}}$ and $X_{\text{answer}}$ to enable the model to use correct tools and answer questions based on the tool’s results.
@@ 3.2 Instruction Tuning
In order to ensure MMedAgent simultaneously performs as both action planner and results aggregator, we adopt the unified dialogue format proposed by Liu et al. (2023a), illustrated in Figure 2. Specifically, upon receiving a user’s input, MMedAgent generates three components in its outputs: (1) Thought, which determines whether MMedAgent can independently solve the user’s instructions or if external tools are required, and if so, identifies the appropriate tool; (2) Action, which enumerate a list of API calls necessary to execute the thought. This comprises two sub-fields: API Name and API Params. If the action list is null, no API call is initiated. (3) Value, which provides a natural language response aggregated by the MLLM along with the outputs from the involved tools. As depicted in Appendix Figure 5, we construct the instruction data by querying GPT-4o through one-shot learning, presenting an example that demonstrates the input and output of MMedAgent. We set a fixed System instruction prompt for each tool and select several examples as conversation templates (User_1 and Assistant_1 in Appendix Figure 5). The tool processes the to generate the instruction data from the dialogue.
@@ 3.3 Medical Tasks and Tools
Our MMedAgent possesses the capability to access a diverse array of tools with the scalability to handle various tasks. As shown in Table 1, we integrate six tools that encompass seven representative tasks in medical domains, i.e., (1) grounding, (2) segmentation with bounding-box prompts (Segmentation), (3) segmentation with text prompts (G-Seg), (4) medical imaging classification, (5) Medical Report Generation (MRG), (6) retrieval augmented generation (RAG), and (7) VQA. Note that no additional tools are required for the VQA task since we utilize LLaVA-Med, which originally supports it. Each tool functions as a specialist, exhibiting exceptional proficiency in executing a specific task across various medical imaging modalities.
@@ 3.3.1 Grounding
Grounding, also known as detection, aims to identify and localize specific objects within an input image by generating the coordinates of bounding boxes containing the objects. To the best of our knowledge, no existing medical models can simultaneously process images from different modalities. Consequently, we propose a generalized grounding tool tailored for the medical domain. Specifically, we choose to fine-tune Grounding DINO Liu et al. (2023b), an open-set object detector, to the medical imaging field.
@@ 3.3.2 Other Tasks
Segmentation involves identifying and delineating the region of interest (ROIs) of an image. In our scenario, we consider interactive segmentation when a bounding box that covers the ROIs is provided. This setting has become popular since the development of Segment Anything (SAM) Kirillov et al. (2023). We select MedSAM Ma et al. (2024a), which fine-tunes SAM to the medical field, as our tool. The prompts are limited to bounding boxes because they provide more precise guidance to SAM Mazurowski et al. (2023). Specifically, in this scenario, we consider the users to provide the position of the bounding box in which MedSAM can be directly applied to obtain the ROI masks.
G-Seg refers to combining grounding with SAM. It aims to address a more common scenario when users specify only a particular object to segment in an image. In this case, we first activate the fine-tuned grounding tool to localize the referred object and then provide its location, in box format, to MedSAM.
Classification aims to identify the most appropriate category for a medical image within a closed set. Specifically, we define a closed set of labels $L$, including organ types, common image modalities, and complex modalities such as ultrasound imaging, hematoxylin, and eosin histopathology. The details of the set $L$ are shown in Appendix A.1. We adopt BiomedCLIP Zhang et al. (2024), which exhibits superior performance in zero-shot and fine-grained classification. The image is classified based on the cosine similarity between the image embedding and each text embedding.
@@
RadFM - - - 25.00 68.13 - 45.38
LLaVA-Med 51.78 65.48 68.58 53.46 70.10 30.44 60.68
LLaVA-Med (Tool in Test) 45.32 52.77 67.91 57.53 74.34 67.55 65.31
MMedAgent (ours) 97.50 102.29 125.89 81.11 121.49 85.55 109.48
Table 2: Performance comparison between MMedAgent and other baselines. RadFM cannot handle grounding and retrieval augmented generation tasks, filled out by “-”. LLaVA-Med refers to the 60K-IM version with only the initial query $X_{q}$ and image $I_{q}$ as input, while LLaVA-Med (Tool in Test) takes $X_{q}$, $I_{q}$ and also the internal output from tools $X_{\text{result}}$ as input.

### v2

# Zhe Xu2,5,*, Shilong Liu6, Haoyu Dong7, Zihao Lin3, Yixin Wang1

@@ Abstract
Multi-Modal Large Language Models (MLLMs), despite being successful, exhibit limited generality and often fall short when compared to specialized models. Recently, LLM-based agents have been developed to address these challenges by selecting appropriate specialized models as tools based on user inputs. However, such advancements have not been extensively explored within the medical domain. To bridge this gap, this paper introduces the first agent explicitly designed for the medical field, named Multi-modal Medical Agent (MMedAgent). We curate an instruction-tuning dataset comprising six medical tools solving seven tasks across five modalities, enabling the agent to choose the most suitable tools for a given task. Comprehensive experiments demonstrate that MMedAgent achieves superior performance across a variety of medical tasks compared to state-of-the-art open-source methods and even the closed-source model, GPT-4o. Furthermore, MMedAgent exhibits efficiency in updating and integrating new medical tools.
@@
Binxu Li1, Tiankai Yan1, Yuanting Pan1, Jie Luo2, Ruiyang Ji3, Jiayuan Ding4, Zhe Xu2,5,*, Shilong Liu6, Haoyu Dong7, Zihao Lin3, Yixin Wang1thanks: Corresponding authors 1Stanford University,  2Harvard Medical School, 3Virginia Tech, 4MSU, 5CUHK, 6Tsinghua University, 7Duke University {andy0207,yxinwang}@stanford.edu, haoyu.dong151@duke.edu, zihaol@vt.edu
@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) have made considerable progress across diverse tasks with inputs from different medical imaging modalities (e.g., Magnetic Resonance Imaging, Computed Tomography, X-ray) in healthcare, including Visual Question Answering (VQA) Moor et al. (2023a); Zhang et al. (2023a); Li et al. (2023), image segmentation Ma et al. (2024a), and Medical Report Generation (MRG) Thawkar et al. (2023); Hamamci et al. (2024), etc. Despite these advancements, MLLMs often exhibit limitations in seamlessly solving multiple tasks across different medical imaging modalities. Although recent large medical models Zhang et al. (2023b); Tu et al. (2024); Wu et al. (2023); Yang et al. (2024); Zhao et al. (2024a) have attempted to address this challenge, they remain limited to handling a narrow range of tasks across a restricted set of imaging modalities and cannot be efficiently extended to new tasks or more imaging modalities. Furthermore, these generalists typically do not provide expert-level responses comparable to those of specialized MLLMs customized for specific tasks.
@@
One way to address this issue is to build an AI Agent, an AI system driven by Large Language Models (LLMs) that integrates various domain expert models as tools. Such a system can understand user instructions, make decisions, and select the appropriate tools to execute any specific task, thereby generating expert-level responses for any given request Xie et al. (2024); Chen et al. (2023); Wang et al. (2024); Liu et al. (2023b); Tao et al. (2023). Despite the significant success of AI agents in the general image domain Tao et al. (2023); Qin et al. (2023); Wang et al. (2023a), there are currently few AI agents developed specifically for the medical domain. Although several works Tang et al. (2023); Schmidgall et al. (2024); Li et al. (2024); Fan et al. (2024) in the medical field use the term “agent” in their methods, they focus on utilizing LLMs to play various roles and collaborate on complex tasks, in which an “agent” refers to a specific role. PathAsst Sun et al. (2024) integrates tool utilization into their framework but specifically designed for pathology tasks.
@@
In this work, we aim to build the first AI agent specifically for the medical domain, termed as Multi-modal Medical Agent (MMedAgent). We choose LLaVA-Med Li et al. (2023) as the backbone and aim to extend its capability to handle various language and multi-modal tasks, including grounding, segmentation, classification, MRG, and Retrieval-Augmented Generation (RAG). These tasks encompass multiple medical imaging modalities, such as MRI, CT, and X-ray, allowing MMedAgent to support a wide range of data types typically encountered in clinical practice. The first step to building MMedAgent is to collect the state-of-the-art (SOTA) methods for each task, hereafter referred to as “tools”. During this phase, we identify a lack of an effective tool for the grounding task, prompting us to fine-tune Grounding DINO Liu et al. (2023c) specifically for medical applications. Next, we build an instruction-based dataset that teaches the agent to select the proper tool(s) when encountering a user instruction and aggregate the outputs from tools to reply to users precisely and comprehensively. The core of our approach involves an end-to-end training regimen through visual instruction tuning Liu et al. (2023b). MMedAgent has demonstrated promising results in various aspects. When evaluated on several complex medical tasks, MMedAgent significantly outperforms sevaral open-source SOTA methods, including LLaVA Liu et al. (2023a), Flamingo-Med Moor et al. (2023b), Yi-VL-34B AI et al. (2024), Qwen-VL-Chat Bai et al. (2023), LLaVA-Med Li et al. (2023) and RadFM Wu et al. (2023), and even surpasses close-source method, GPT-4o OpenAI (2024), on average. It also enhances MMedAgent’s backbone, i.e., LLaVA-Med, original capability in the VQA task, as well as exhibits efficient capability in learning new tools. Our code and web UI are available at https://github.com/Wangyixinxin/MMedAgent, and a demonstration of the user interface is provided in Appendix Figure 6.
@@
  • We build the first open-source instruction tuning dataset for general-purpose multi-modal medical agents.
@@ 2.1 Medical MLLMs
Recent progress on LLMs has been made on multi-modal conversational capability Moor et al. (2023a); Zhang et al. (2023b); Tu et al. (2024); Zhang et al. (2023c, a); Thawkar et al. (2023); Sun et al. (2024); Wu et al. (2023); Li et al. (2023); Ma et al. (2024a); Yang et al. (2024); Zhao et al. (2024a); Hamamci et al. (2024). Owing to the diversity inherent in medical data and tasks, LLMs have initially been localized to specific imaging domains such as X-ray Thawkar et al. (2023), CT Hamamci et al. (2024), and histology Sun et al. (2024), or tailored for different tasks such as segmentation Ma et al. (2024a); Lei et al. (2023) and medical report generation Wu et al. (2023). In contrast, generalist models expand these capabilities by enabling a single LLM to cover a wider range of imaging modalities and tasks by enlarging the pre-training datasets greatly Zhang et al. (2023b); Li et al. (2023); Zhao et al. (2024a); Liu et al. (2023b); Yang et al. (2024). Although generalist models are capable of handling a wide range of medical modalities and tasks, they face limitations in scalability when incorporating additional skills and lack specialization in specific tasks.
@@ 2.2 AI Agent
A multi-modal AI Agent is a system that achieves users’ general-purpose goals by perceiving the environment and making decisions based on the perceptions Xie et al. (2024); Wooldridge and Jennings (1995). Recent works utilize LLMs as planners to understand multi-modal input from environments and make decisions to call different tools to achieve goals. Based on whether the LLM is open source or not, (Xie et al., 2024) classifies multi-modal AI Agents into two types: (i) closed-source LLMs as planners, which utilize prompt technique to enable LLMs to make decisions Chen et al. (2023); Wang et al. (2024); (ii) fine-tuned LLMs as planners, where an LLM is fine-tuned to understand instructions, make decisions, and call tools/APIs Liu et al. (2023b); Tao et al. (2023); Zhang et al. (2024b). MMedAgent belongs to the second type.
Multi-modal AI Agents have achieved great success in various applications. For example, (Tao et al., 2023; Gur et al., 2023; Zhan and Zhang, 2023) apply agents to control the website or user interface. Some works Qin et al. (2023); Wang et al. (2023c) focus on robotics or embodied AI which applies multi-modal LLMs to perceive and interact with real environments. Most works concentrate on multi-modal understanding, or generation, especially image, video, or audio Liu et al. (2023b); Wang et al. (2023a); Zhang et al. (2023e). However, these works are limited to the natural domains. To the best of our knowledge, we are the first to build a more versatile medical AI Agent, which encompasses a broader spectrum of image modalities, including MRI, CT, X-ray, and histology.
@@ 3.1 Workflow
Following LLaVA-Plus Liu et al. (2023b), the objective of MMedAgent is to learn to utilize a wide range of multi-modal medical tools, extending the MLLMs’ capabilities to analyze and accomplish various medical tasks. As shown in Figure 1, the workflow consists of four parts: (1) users provide an instruction $X_{q}$ and a medical image $I_{q}$; (2) MLLM works as an action planner, which understands $X_{q}$ and $I_{q}$ and then generates a formatted instruction $X_{\text{tool}}$ to call a specific tool. (3) The tool is executed given $I_{q}$ and the output $X_{\text{result}}$ of the tool is sent to the MLLM. (4) The MLLM aggregates the output with $X_{q}$ and $I_{q}$ and generates the final answer $X_{\text{answer}}$ to users. We train the agent end-to-end with an auto-regressive objective on the generated sequence, i.e., $X_{\text{tool}}$ and $X_{\text{answer}}$, to enable the model to use correct tools and answer questions based on the tool’s results.
@@ 3.2 Instruction Tuning
In order to ensure MMedAgent simultaneously performs as both action planner and results aggregator, we adopt the unified dialogue format proposed by Liu et al. (2023b), illustrated in Figure 2. Specifically, upon receiving a user’s input, MMedAgent generates three components in its outputs: (1) Thoughts, which determine whether MMedAgent can independently solve the user’s instructions or if external tools are required, and if so, identifies the appropriate tool; (2) Actions, which enumerate a list of API calls necessary to execute the thought. This comprises two sub-fields: API Name and API Params. If the action list is null, no API call is initiated. (3) Value, which provides a natural language response from the MLLM. During the first round, it indicates the chosen tool(s); during the second round, it represents the final output that addresses the user’s initial request. As depicted in Appendix Figure 5, we construct the instruction data by querying GPT-4o through one-shot learning, presenting an example that demonstrates the input and output of MMedAgent. We set a fixed System instruction prompt for each tool and select several examples as conversation templates (User_1 and Assistant_1 in Appendix Figure 5). The tool processes the generation of instruction data from the dialogue.
@@ 3.3 Medical Tasks and Tools
Our MMedAgent possesses the capability to access a diverse array of tools with the scalability to handle various tasks. As shown in Table 1, we integrate six tools that encompass seven representative tasks in medical domains, i.e., (1) grounding, (2) segmentation with bounding-box prompts (B-Seg), (3) segmentation with text prompts (G-Seg), (4) medical imaging classification, (5) Medical Report Generation (MRG), (6) retrieval augmented generation (RAG), and (7) VQA. Note that no additional tools are required for the VQA task since we utilize LLaVA-Med, which originally supports the task, as the backbone. Each tool functions as a specialist, exhibiting exceptional proficiency in executing a specific task across various medical imaging modalities.
@@ 3.3.1 Grounding
Grounding, also known as detection, aims to identify and localize specific objects within an input image by generating the coordinates of bounding boxes containing the objects. To the best of our knowledge, no existing medical models can simultaneously process images from different modalities. Consequently, we propose a generalized grounding tool tailored for the medical domain. Specifically, we choose to fine-tune Grounding DINO Liu et al. (2023c), an open-set object detector, to the medical imaging field.
@@ 3.3.2 Other Tasks
B-Seg involves identifying and delineating the region of interest (ROIs) of an image when a bounding box that covers the ROIs is provided. It is a type of interactive segmentation, which has become popular since the development of Segment Anything (SAM) Kirillov et al. (2023). We select MedSAM Ma et al. (2024a), which fine-tunes SAM to the medical field, as our tool. The prompts are limited to bounding boxes because they provide more precise guidance to SAM Mazurowski et al. (2023). Specifically, in this scenario, we consider the users to provide the position of the bounding box in which MedSAM can be directly applied to obtain the ROI masks.
G-Seg refers to combining grounding with SAM. It aims to address a more common scenario when users specify only a particular object to segment in an image. We first activate the fine-tuned grounding tool to localize the referred object and then provide its location, in box format, to MedSAM. Note that this task also presents MMedAgent’s capability to address complex tasks with more than one tools, an ability not presented in previous works.
Classification aims to identify the most appropriate category for a medical image within a closed set. Specifically, we define a closed set of labels $L$, including organ types, common image modalities, and complex modalities such as ultrasound imaging, hematoxylin, and eosin histopathology. The details of the set $L$ are shown in Appendix A.1. We adopt BiomedCLIP Zhang et al. (2024a), which exhibits superior performance in zero-shot and fine-grained classification. The image is classified based on the cosine similarity between the image embedding and each text embedding.
@@
Flamingo-Med Moor et al. (2023b) 13.11 15.87 15.33 23.56 16.59 - 14.68 1.16
RadFM Wu et al. (2023) - - - 25.00 68.13 - 45.38 3.59
LLaVA-Med Li et al. (2023) 51.78 65.48 68.58 53.46 70.10 30.44 60.68 4.80
Yi-VL-34B AI et al. (2024) 63.23 79.40 68.32 76.02 72.95 14.67 64.08 5.07
LLaVA-Med (Tool in Test) 45.32 52.77 67.91 57.53 74.34 67.55 65.31 5.17
Qwen-VL-Chat Bai et al. (2023) 61.34 65.90 62.38 88.40 73.41 78.80 76.21 6.03
LLaVA-34B Liu et al. (2023a) 76.75 84.85 80.75 96.04 80.27 91.64 86.52 6.84
MMedAgent (ours) 97.50 102.29 125.89 81.11 121.49 85.55 109.48 8.66
Table 2: Performance comparison between MMedAgent and other baselines. Cls. stands for classification. We report the relative scores for all tasks and the absolute (abs.) scores for overall performance in the last column. “-” indicates the tasks that the corresponding model is not applicable to. LLaVA-Med refers to the 60K-IM version with only the initial query $X_{q}$ and image $I_{q}$ as input, while LLaVA-Med (Tool in Test) takes $X_{q}$, $I_{q}$ and also the internal output from tools $X_{\text{result}}$ as input.